-21 -2 1 1 Let A Determine whether the system Ac = b is consistent for every be R3. --2)
Answer #1

2 L A = -2 1 -2 arank Here first we calculate -2 1 Interchange Root and 2 Rood 2 1 A 1 1 -2 Le = Rq + 2h Rg = R3-R R2 = 1/R2